Grade 12Physical Chemistry

How much copper can be obtained from 100 g of copper sulphate (CuSO4)

sudhanshu
12 Years agoGrade 12

3 Answers

Gaurav
11 Years ago
Hello Student
Molar mass of copper sulphate = 159.609g/mol
100 g Copper sulphate = 100/159.6 = 0.63 mole of Cu
1 Mole of Cu = 63.5 g
0.63 mole of Cu = 0.39 g
Rishabh Sharma
11 Years ago
CuSO4  → CuO + SO3
 
So, 1 mol  of CuSO4 reacts with 1 mol of CuO.
1 mol of CuO contains 1 mol of Cu.
therefore calulate the moles of CuSO4 and equal moles will be of Cu.
Prakash
8 Years ago
CuSO4 ~ Cu+So4^2-Hence 1 mole of CuSO4 provides 1 mole of Cu159.5g of CuSO4 gives to 63.5g of CuMolar mass of CuSO4 is 159.5g and Cu is 63.5g1g of CuSO4 = 63.5/159.5 g of cu100 g of CuSO4 =0.398×100 g of CuHence 100g of CuSO4 will give 39.8 g of copper